OpenStudy (anonymous):

4x+y=16 2x+3y=-2 (5,-4) (-5,4) (4,-5) (5,4)

OpenStudy (anonymous):

Name equations 4x+y=16 i) 2x+3y=-2 ii) ii)x2 to get 4x+6y=-4 iii) equation iii)-ii) to get y sub y into i) ii) or iii) to get x

OpenStudy (anonymous):

i mean equation iii)-i) to get y sub y into i) ii) or iii) to get x

OpenStudy (anonymous):

to get y=-4 x=5